I want to see the steps on solving this problem; 6((2y/3))-3y+2=0.
Thank you

Answer by Cromlix(4381)

6((2y/3))-3y+2=0.
In the brackets the external 6
is cancelled down to 2 by the 3
below the 2y.
3(2y) - 3y + 2 = 0
6y - 3y + 2 = 0
Collect like terms
3y + 2 = 0
3y = -2
y = -2/3
Hope this helps.
:-)